OpenStudy (anonymous):

i have a probability question . two dices are tossed once. let the random variable be t he sum of the up faces on the dice. A). find and graph the probability distribution of the random variable. and b) calculate the mean (or expectation) of this distribution

OpenStudy (dumbcow):

there are 36 possibilities so prob = 1/36 sum can be 2,3...,12 1 way to get 2 and 12 --> p = 1/36 2 ways to get 3 and 11 -->p = 2/36 3 ways to get 4 and 10 --> p=3/36 4 ways to get 5 and 9 --> p = 4/36 5 ways to get 6 and 8 --> p=5/36 6 ways to get 7 --> p =6/36 E = 7